OpenStudy (anonymous):

So I'm given this: f(X)={(3,5), (2,4), (1,7)} g(X)=sqrt{x-3} h(x)= {(3,2), (4,30, (1,6)} k(x)= x^2+5 And then asked: (f+h)(1)=? what do I do?

OpenStudy (anonymous):

find \(f(1)\) then find \(h(1)\) and then add

OpenStudy (anonymous):

do you know what \(f(1)\) is? "no" is a fine answer, i am just asking if you don't know i will show you what it is

OpenStudy (anonymous):

i was thinking f(h(1))

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

it is not asking for \(f(h(1))\) it is asking for \(f(1)+h(1)\)

OpenStudy (anonymous):

\[f(x)=\{(3,5), (2,4), (1,7)\}\] to find \(f(1)\) find the ordered pair where the first coordinate is \(1\) then \(f(1)\) is the second coordinate of that pair

OpenStudy (anonymous):

nothing to compute here, we see \((1,7)\) and know that \(f(1)=7\)

OpenStudy (anonymous):

\[h(x)= \{(3,2), (4,30, (1,6)\}\] what is \(h(1)\)?

OpenStudy (anonymous):

6!

OpenStudy (anonymous):

exactly!

OpenStudy (anonymous):

so what you are being asked is really "what is \(7+6\) ? "

OpenStudy (anonymous):

awesome, I understand. how about (f (a circle symbol) h) 3

OpenStudy (anonymous):

\[f\circ h(x)\]?

OpenStudy (anonymous):

yes! do i mulitply the 3?

OpenStudy (anonymous):

oh i see \(f\circ h(3)\)

OpenStudy (anonymous):

oh hell no

OpenStudy (anonymous):

that is not a multiplication that means \[f(h(3))\]

OpenStudy (anonymous):

work always from the inside out the first thing you need to find it \(h(3)\)

OpenStudy (anonymous):

\[h(x)= \{(3,2), (4,30, (1,6)\}\] what is \(h(3)\) ?

OpenStudy (anonymous):

(3,2)

OpenStudy (anonymous):

\(h(3)\) is a number, not an ordered pair \((3,2)\) is an ordered pair

OpenStudy (anonymous):

the first coordinate is \(3\) and the second coordinate is \(h(3)\)

OpenStudy (anonymous):

recall that you said \(h(1)=6\) right? it was a number, not \((1,6)\)

OpenStudy (anonymous):

so what is \(h(3)\) ?

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

the ordered pair is \((3,2)\) so \(h(3)=2\) we are not done yet though

OpenStudy (anonymous):

\[f(h(3))=f(2)\] since we know \(h(3)=2\) now you have to find \(f(2)\)

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

\[f(x)=\{(3,5), (2,4), (1,7)\}\] yes

OpenStudy (anonymous):

that is your answer lets put it together all in one line

OpenStudy (anonymous):

\[f\circ h(3)=f(h(3))=f(2)=4\]

OpenStudy (anonymous):

this is the last questions ill ask but what if theres a negative one exponent? f^-1(x)=?

OpenStudy (anonymous):

\[f^{-1}(x)\] is the "inverse function" of \(f\) if your function is given by ordered pairs, all you have to do is switch the coordinates of each ordered pair i.e. put the second one first and the first one second

OpenStudy (anonymous):

is it clear what i mean by that?

OpenStudy (anonymous):

oh I see, and if i have K^-1(x)= is it y^2+5

OpenStudy (anonymous):

let me check

OpenStudy (anonymous):

ok for \(f^{-1}\) since \[f(x)=\{(3,5), (2,4), (1,7)\}\] you have \[f^{-1}(x)=\{(5,3), (4,2), (7,1)\}\] see what i did for that?

OpenStudy (anonymous):

yes thats what i got, but does that apply to k^-1(X)? do i change X^2 to its inverse y^2? or its there more to it?

OpenStudy (anonymous):

you can write \[k(x)= x^2+5\]replace it by \[y=x^2+5\] then switch \(x\) and \(y\) because that is what the inverse function does, and write \[x=y^2+5\] then solve for \(y\) in two steps

OpenStudy (anonymous):

\[x=y^2+5\] \[x-5=y^2\] \[\pm\sqrt{x-5}=y\]

OpenStudy (anonymous):

as you can see the inverse of \(k\) is not a function, because of that \(\pm\) out front the reason the inverse is not a function is because \(k(x)=x^2+5\) is not "one to one" for example \(k(2)=9\) and \(k(-2)=9\)as well

OpenStudy (anonymous):

so the inverse is not a function because \((2,9)\) and \((-2,9)\) are both on the graph of \(k\) so the graph of the inverse would have both \((9,2)\) and \((9,-2)\) a function cannot have that

OpenStudy (anonymous):

thank you so much for your help!

OpenStudy (anonymous):

yw